# Toward High-Performance LLM Serving: A Simulation-Based Approach for Identifying Optimal Parallelism

### v1

@@ Abstract.
Large Language Models (LLMs) have achieved significant success in various domains, and serving LLMs efficiently has become crucial. LLMs are often served with multiple devices using parallelism techniques like data, pipeline, and tensor parallelisms. Each parallelism presents trade-offs between computation, memory, and communication overhead, making it challenging to determine the optimal parallel execution plan. Moreover, input workloads also impact parallelism strategies. Tasks with long prompts like article summarization are compute-intensive, while tasks with long generation lengths like code generation are often memory-intensive; these differing characteristics result in distinct optimal execution plans. Since searching for the optimal plan via actual deployment is prohibitively expensive, we propose APEX, an LLM serving system simulator that efficiently identifies an optimal parallel execution plan. APEX performs dynamism-aware simulation, which captures the complex characteristics of iteration-level batching, a technique widely used in state-of-the-art LLM serving systems. APEX leverages the repetitive structure of LLMs to reduce design space, maintaining a similar simulation overhead, even when scaling to trillion scale models. APEX supports a wide range of LLMs, device clusters, etc., and it can be easily extended to new models or clusters through its high-level templates. We run APEX simulations using a CPU and evaluate the optimal parallel execution plans found by APEX using a cluster with 8 H100 GPUs. The simulations cover a myriad of LLM architectures, precision formats, cluster configurations, and input workloads. We show that APEX can find optimal execution plans that are up to 4.42$\times$ faster than heuristic plans in terms of end-to-end serving latency. APEX also reports a set of metrics used in LLM serving systems, such as time per output token (TPOT) and time to first token (TTFT). Furthermore, APEX can identify an optimal parallel execution plan within 15 minutes using a CPU. This is 71$\times$ faster and 1234$\times$ more cost-effective than actual deployment on a GPU cluster using cloud services. APEX will be open-sourced upon acceptance. Finally, we show that LLM service providers can utilize APEX to meet service-level objectives and explore hardware design space to build high-performance LLM serving platforms.
@@ 1. Introduction
Large Language Models (LLMs) have been successfully applied to various applications, such as code generation (Fakhoury et al., 2024; Zhong and Wang, 2024), question-answering (Zhuang et al., 2023; Lucas et al., 2024), and many more (Zhong et al., 2023; Wang et al., 2024; Zhou et al., 2024). Serving an LLM is both memory- and compute-intensive; therefore, leveraging a cluster with multiple devices is essential to gain computing power and memory resources for achieving high performance. Several techniques have been proposed to parallelize LLM on multiple devices, such as data parallelism (DP) (Li et al., 2020b), pipeline parallelism (PP) (Huang et al., 2019), tensor parallelism (TP) (Shoeybi et al., 2020), etc. Each parallelism has its pros and cons. Techniques like TP are more memory-efficient than DP as it does not produce model replicas, but they incur expensive collective communication overhead (Shoeybi et al., 2020). On the other hand, PP has lower communication overhead but can suffer from workload imbalance (Huang et al., 2019). To balance the trade-offs, parallelism techniques can be adopted in a hybrid manner, potentially leading to better serving performance than relying on a single type of parallelism. However, determining the optimal parallel execution plan is challenging due to the many factors involved, including computation workload, network traffic, memory efficiency, etc. Furthermore, the optimal parallel execution plan also depends on the characteristics of the input requests. Some requests lead to long prompts with short generation lengths, such as summarizing an article, while some lead to short prompts and long generation lengths, like code generation and storytelling. The former type of request is compute-intensive, while the latter is memory-intensive. As a result, these two types of requests favor different parallelism techniques to optimize performance. A common practice is to adopt heuristic execution plans instead, such as applying TP within the same node and PP across different nodes (Miao et al., 2024; vLLM, v060). Yet, a recent study has shown that applying such heuristics approaches can be up to 2$\times$ slower than the optimal configuration (Isaev et al., 2023); this calls for the need for service providers to search for and adopt an optimal parallel execution plan rather than relying on heuristics.
@@
To find an optimal execution plan, a straightforward approach is to deploy and evaluate various parallel execution plans directly; however, this approach is prohibitively expensive, as it could take thousands of GPU hours to assess multiple execution plans (Agrawal et al., 2024b, a). Note that this high searching cost cannot be amortized, as the optimal configuration varies depending on the model and the workload characteristics of the input requests. An alternative is to search via modeling-based approaches, such as developing performance models (Isaev et al., 2023) or building simulators (Yu et al., 2021; Duan et al., 2024) to estimate the performance of a given parallel execution plan; nevertheless, existing solutions developed for Deep Neural Network (Yu et al., 2021; Duan et al., 2024) or LLM training (Isaev et al., 2023; Bang et al., 2024) cannot be applied to LLM serving, as such systems introduce unique challenges: (1) Dynamism of iteration-level batching: Unlike conventional ML systems that adopt static batching, which waits for all requests in the current batch to be completed before batching new requests, state-of-the-art LLM serving systems (Kwon et al., 2023; Agrawal et al., 2024b; Zhong et al., 2024) adopt iteration-level batching (Yu et al., 2022) to achieve high serving performance. With iteration-level batching, incoming requests are continuously added to the processing batch whenever memory becomes available. This makes it challenging to model the system, as the batch size changes dynamically during serving. Additionally, some requests within the batch may be in the prefill stage, while others are in the generation stage. This interleaving of stages further complicates the modeling, as these stages have significantly different computational characteristics (details in Section 2.1). (2) Exponentially-growing design space: Although modeling-based approaches provide a potentially time-efficient solution to search for the optimal parallel execution plan, they can still incur substantial overhead as the design space grows exponentially with respect to the model size and the number of devices in the cluster. Since LLM serving (Kwon et al., 2023; Agrawal et al., 2024b; Zhong et al., 2024) often involves large models and clusters, it is non-trivial to develop a solution that can search for an optimal parallel execution plan within a reasonable timeframe. (3) Adapting for continuously-evolving systems: LLM serving system is an emerging area, with continuous evolution in model architecture, hardware platforms, and system optimizations such as quantization (Lin et al., 2024; Hooper et al., 2024) and parallelism techniques (Rajbhandari et al., 2022; Liu et al., 2023). Consequently, a performance model or a simulator can easily become obsolete and inapplicable to state-of-the-art LLM serving systems. Thus, keeping pace with the rapid evolution of such systems is necessary but also challenging.
To this end, we propose APEX, an extensible and dynamism-aware simulator for automated parallel execution in LLM serving. APEX takes an LLM, a set of input requests with various context and generation lengths, and a device cluster as inputs, and generates an optimal parallel execution plan for LLM serving. Specifically, for a given LLM model and device cluster, APEX first generates various parallel execution plans, each representing a unique way to parallelize the model by combining various parallelism techniques, such as tensor, data, pipeline, and expert parallelisms. APEX then evaluates each plan by estimating the execution time of serving the input requests through simulation. APEX performs dynamism-aware simulation that is capable of modeling the complex characteristics of iteration-level batching, which involves concurrently serving requests of varying lengths and stages (i.e., prefill and generation stage). After the simulation, APEX provides a comprehensive evaluation for each parallel execution plan, which includes multiple metrics used in LLM serving systems, such as time per output token (TPOT), time to first token (TTFT), end-to-end serving latency, P95 latency, among others (Miao et al., 2023). A parallel execution plan with the lowest end-to-end serving latency is chosen as the optimal plan. Despite performing complex dynamism-aware simulations, APEX remains time-efficient by leveraging the repetitive structure of transformer layers in LLMs, which significantly reduces the design space; this allows APEX to scale to trillion-scale models on multi-node device clusters. APEX supports a broad range of LLMs, device clusters, quantization formats, and parallelism techniques, which are essential for modeling state-of-the-art LLM serving systems. Recognizing the rapid evolution of LLM serving systems, APEX is designed to be modular and extensible. We capture the high-level abstractions of LLMs and device clusters, and develop software templates based on these abstractions. This approach allows new models and device clusters to be easily supported with minimal coding effort in the templates. Additionally, APEX leverages operation-level profiling results to estimate the execution time of parallel execution plans on a given device cluster. Profiling-based estimation enables APEX to adapt to different clusters by collecting profiling data from the underlying platform. Note that the device cluster APEX supports is not limited to GPU clusters, but also AI-accelerator clusters such as TPU (Jouppi et al., 2023), Intel Gaudi (Intel, 2024), etc. We plan to open-source APEX upon paper acceptance. We further demonstrate that, in addition to identifying optimal parallel execution plans, APEX can also assist LLM service providers in meeting service-level objectives (SLOs) and provide valuable insights for building LLM-serving hardware platforms. The key contributions of this work are:
@@
  • We propose APEX, a dynamism-aware simulator that captures the complex characteristics of LLM serving. APEX automatically finds an optimal parallel execution plan for a given LLM, a device cluster, and a set of input requests.
  • APEX supports a wide range of LLMs, parallelism techniques, quantization formats, and device clusters. APEX can also be easily extended to new models and device clusters.
  • We evaluate APEX using four LLMs on three datasets with distinct workloads. APEX identifies optimal parallel execution plans that are up to 4.42$\times$ faster than heuristic plans in terms of the end-to-end serving latency.
  • APEX provides high-fidelity simulation, achieving less than 10% relative error on average when predicting the speedup of optimal execution plans over the baseline plans.
  • APEX is time-efficient and cost-effective, capable of finding an optimal parallel execution plan within 15 minutes on a CPU; this is 71$\times$ faster and 1234$\times$ cost-effective than actual deployment on a GPU cluster using cloud services.
  • APEX is highly scalable, maintaining similar simulation overhead when scaling from billion-scale to trillion-scale models.
@@
  • APEX features comprehensive evaluation, which can help service providers meet SLOs. APEX also provides insightful suggestions for building LLM-serving hardware platforms.
@@ 2.1. LLM Inference
Large Language Models (LLMs) are a type of ML models that are built upon transformer architecture (Vaswani et al., 2017). LLM inference takes a user prompt as input, and generates a response, token by token, sequentially. The sequential token generation process is also known as the auto-regressive generation, which utilizes the previously generated tokens as input to predict the next token. The inference process of LLMs consists of two stages:
Prefill Stage: In the prefill stage, LLM processes the input request (i.e., prompt) to set up intermediate states (keys and values) that are used to predict the first token. Unlike token generation, the computation of prefill does not rely on previously generated output tokens, allowing the tokens of the input request to be processed in parallel all at once. This high degree of parallelism makes the prefill stage compute-bound (Zhong et al., 2024).
Generation Stage: During the generation phase, LLM generates output tokens autoregressively until a stopping criterion is met. The generation of each token depends on all previous tokens’ output states (keys and values). The generation phase is often memory-bound (Lin et al., 2024), as the latency mainly depends on the speed of data transfer for the output states from the memory, rather than on the computation itself.
@@ 2.2. LLM Serving and Iteration-Level Batching
LLMs are often hosted by service providers in the cloud (Li et al., 2024). Users send requests to these providers through APIs or chatbots and then receive responses generated by the LLMs. While a single LLM inference processes a single input request and generates one output response, LLM serving concurrently processes multiple input requests from the users and generates multiple output responses, i.e., multiple LLM inferences happen in parallel. For LLM serving, it is critical to achieve both low latency and high throughput: low latency is necessary to meet service-level objectives, while high throughput enables service providers to serve a large number of users simultaneously. Due to the autoregressive nature of LLMs, the lengths of the generated responses can vary significantly. Consequently, the commonly used static batching leads to suboptimal performance in LLM serving, as all the batched requests need to wait for the longest response to be generated to proceed to process new requests. To overcome this inefficiency, iteration-level batching (Yu et al., 2022) is proposed. Iteration-level batching continuously schedules newly arrived request(s) into the existing batch whenever GPU memory becomes available, rather than waiting for all the requests to be completed. Iteration-level batching significantly improves the serving throughput and is widely adopted in state-of-the-art LLM serving systems (Kwon et al., 2023; Agrawal et al., 2024b; Zhong et al., 2024; NVIDIA, 2024).
2.3. Device Clusters and Interconnections
Serving an LLM is both memory- and compute-intensive. To achieve high serving throughput and accommodate the large model size, LLMs are often deployed on a cluster of devices. The devices are connected in a tree-based network topology, which is one of the most popular topologies for node deployment. The memory bandwidth and latency is uniform within each level. Figure 1 illustrates an example of a commonly used two-level device cluster. Devices within the same node are typically connected by high-bandwidth interconnects like NVLink (Li et al., 2020a) at level 1, while devices across different nodes are connected via inter-node networks such as InfiniBand (Liu and Guitart, 2022) at level 2. In addition to GPUs, clusters of AI accelerators are also emerging, such as TPU clusters (Jouppi et al., 2023) and Gaudi clusters (Intel, 2024). These clusters also utilize tree-based network topologies for interconnection and can be abstracted similarly as GPU clusters (Choquette, 2023).
Figure 1. An example of a two-level device cluster hierarchy, where leaves represent the devices. The memory bandwidth and latency are uniform within the same level.
Figure 2. System overview of APEX
@@ 2.4. LLM Parallelisms
Data Parallelism (DP): In DP (Li et al., 2020b), the model is replicated across multiple devices. The input requests are split into micro-batches and distributed to each model replica for processing. DP incurs no communication overhead as each micro-batch is processed independently. However, having model replicas incurs large memory overhead, and reduces the number of requests that can be batched concurrently.
@@
Expert Parallelism (EP): EP (Rajbhandari et al., 2022) is a special type of parallelism for Mixture of Experts (MoE) models. EP deploys different experts across various devices. For MoE models, only a subset of experts are activated for each input token. Thus, EP can result in workload imbalance and resource under-utilization if none of the experts on the device are activated; while this can be avoided by adopting TP, where each device is assigned with a partition of all the experts, EP incurs a lower communication overhead than TP.
@@ 3.1. System Overview
We depict the system overview of APEX in Figure 2. Initially, the Offline Profiler (Section 3.2) of APEX takes a set of profiling scripts to obtain the performance information of the underlying platform, and the results are stored for APEX’s profiling-based simulation. Profiling is an offline process that is only performed once when porting to a new cluster with unknown devices (e.g., porting from an A100 to an H100 GPU cluster.) Next, given an LLM and device cluster, the Parallel Execution Plan Generator converts the model into an intermediate representation (Section 3.3) and generates various parallel execution plans (Section 3.4 and 3.5); each plan represents a unique way to map the LLM onto the device cluster. The Iteration-Level Batching Module (Section 3.6) takes the generated parallel execution plans and a set of request traces as inputs, and starts batching the input requests. The Batching Module determines whether a request should be added to or removed from the batch in each iteration, and reports the active requests to the Dynamism-Aware Simulator (Section 3.7). Based on the active requests in the batch, the Dynamism-Aware Simulator estimates the execution time for each iteration and keeps track of multiple LLM system metrics such as time to first token (TTFT), time per output token (TPOT), P95 latency, etc. After all requests have been processed, the Dynamism-Aware Simulator produces a Simulation Report for the corresponding parallel execution plan. This process is repeated for all the generated parallel execution plans, and the plan with the lowest end-to-end latency is selected as the optimal plan.
3.2. Offline Profiler
The Offline Profiler collects performance information of the underlying platform, where the results are used by the Dynamism-Aware Simulator (Section 3.7) to estimate the execution time of different parallel execution plans. The Offline Profiler leverages the fact that LLMs are based on transformer architecture, which consists of similar operations, and performs operation-level profiling, such as measuring the computation time of multi-head attention, which is used in the attention layers, and general matrix multiplication (GEMM), which is used in the feedforward layers. Profiling the key transformer operations allows APEX to support various LLMs without exhaustively profiling each model, as they can be broken down into the same set of operations. In addition to computation operation, the Offline Profiler also profiles the collective communication overheads on the cluster, e.g., the time to perform AllReduce, ReduceScatter, All-to-All. This allows APEX to support various parallelisms, which require different collective communications. For the computation operations, the profiling is performed by going through various sequence lengths, number of attention heads, hidden dimensions, etc. For the collective communication operations, the profiling is performed by varying the data transfer sizes and also the number of devices involved, both within a node and across nodes, in the collective operation. Operation-level profiling produces profiling results that can support various LLMs, request traces, parallelisms, and device clusters of different sizes. The profiling only needs to be performed once when porting to a new cluster with unknown devices, such as porting from an A100 GPU cluster to a H100 GPU cluster. Therefore, for the same type of device cluster, the profiling overhead is a one-time cost that can be amortized.

### v2

# APEX: An Extensible and Dynamism-Aware Simulator for Automated Parallel Execution in LLM Serving

@@ Abstract.
Efficiently serving Large Language Models (LLMs) requires selecting an optimal parallel execution plan, balancing computation, memory, and communication overhead. However, determining the best strategy is challenging due to varying parallelism techniques (data, pipeline, tensor) and workload characteristics (e.g., compute-intensive tasks with long prompts vs. memory-intensive tasks with long generation). We propose APEX, an LLM serving system simulator that efficiently identifies optimal parallel execution plans by considering key factors of LLM serving systems, such as memory usage, batching behavior, etc. APEX performs dynamism-aware simulation to model iteration-level batching, and leverages LLMs’ repetitive structure to reduce design space, scaling efficiently to trillion-scale models. APEX abstracts the key components of LLM serving systems—including the model, batching module, quantization formats, and device clusters—enabling the simulator to be general and extensible. Simulating on a CPU, APEX evaluates execution plans for various device clusters, covering diverse LLMs and workloads. APEX finds plans up to 3.37$\times$ faster than heuristics, and also plans that reduce energy consumption by up to 45% compared to latency-optimal plans. APEX performs comprehensive evaluations, reporting key system metrics like time per output token and time to first token, which can help service providers meet SLOs. APEX identifies an optimal plan within 15 minutes on a CPU, making it 71$\times$ faster and 1234$\times$ more cost-effective than cloud-based GPU deployment. APEX can be accessed at: https://github.com/microsoft/apex_plus
@@ 1. Introduction
Large Language Models (LLMs) have been successfully applied to various applications, such as code generation (Fakhoury et al., 2024; Zhong and Wang, 2024), question-answering (Zhuang et al., 2023; Lucas et al., 2024), and many more (Zhong et al., 2023; Wang et al., 2024; Zhou et al., 2024). Serving an LLM is both memory- and compute-intensive, necessitating multi-device clusters to achieve high performance. Several techniques have been proposed to parallelize LLM, such as data parallelism (DP) (Li et al., 2020b), pipeline parallelism (PP) (Huang et al., 2019), tensor parallelism (TP) (Shoeybi et al., 2020), etc. Each parallelism has its pros and cons: TP is memory efficient, as it does not produce model replicas like DP, but it incurs an expensive communication overhead (Shoeybi et al., 2020). PP reduces communication overhead, but can suffer from workload imbalance (Huang et al., 2019). Hybrid parallelism can effectively balance the trade-offs to improve performance, but identifying the optimal parallel execution plan is non-trivial due to the many factors involved, including computation workload, network traffic, memory efficiency, etc. Furthermore, the optimal parallel execution plan also depends on the characteristics of the input requests. Compute-intensive tasks (e.g., tasks with long prompts like summarization) and memory-intensive tasks (e.g., tasks with long generations like coding) favor different parallelism strategies. A common practice is to adopt heuristic execution plans instead, such as applying TP within the same node and PP across different nodes (Miao et al., 2024; vLLM, v060). Yet, recent studies show that heuristics plans can be 2$\times$ slower than optimal configurations (Isaev et al., 2023), underscoring the need for service providers to search for and adopt an optimal parallel execution plan rather than relying on heuristics.
@@
To find an optimal execution plan, a straightforward way is to deploy and evaluate parallel execution plans exhaustively; however, this is prohibitively expensive as it would take thousands of GPU hours (Agrawal et al., 2024b, a). An alternative is to leverage modeling-based approaches, such as performance models (Isaev et al., 2023) or simulators (Yu et al., 2021; Duan et al., 2024), to estimate the performance of parallel execution plans; nevertheless, existing solutions developed for Deep Neural Network (Yu et al., 2021; Duan et al., 2024) or LLM training (Isaev et al., 2023; Bang et al., 2024) cannot be applied to LLM serving, as serving systems introduce unique challenges: (1) Dynamism of iteration-level batching: Unlike conventional ML systems that use static batching, where all requests in a batch must be completed before new ones are added, LLM serving systems (Kwon et al., 2023; Agrawal et al., 2024b; Zhong et al., 2024) employ iteration-level batching (Yu et al., 2022). In this approach, incoming requests are continuously added to the processing batch as memory becomes available, resulting in dynamically changing batch sizes which are hard to model. Moreover, some requests within the batch may be in the prefill stage, while others are in the decode stage. This interleaving of stages further complicates the modeling, as these stages have significantly different computational characteristics (details in Section 2.1). (2) Exponentially-growing design space: The design space grows exponentially with respect to the model size and the number of devices in the cluster. Since LLM serving often involves large models and clusters, it is non-trivial to develop a solution that can search for an optimal parallel execution plan within a reasonable timeframe. (3) Adapting for continuously-evolving systems: LLM serving system is an emerging area, with continuous evolution in model architecture, hardware platforms, and system optimizations such as quantization (Lin et al., 2024; Hooper et al., 2024) and parallelism techniques (Rajbhandari et al., 2022; Liu et al., 2023). Consequently, a performance model or a simulator can easily become obsolete and inapplicable to the latest LLM serving systems. Thus, keeping pace with the rapid evolution of such systems is necessary but also challenging.
To this end, we propose APEX, an extensible and dynamism-aware simulator for Automated Parallel EXecution in LLM serving. Given an LLM, a device cluster, and input requests with varying context/generation lengths, APEX generates an optimal parallel execution plan for LLM serving. APEX first generates various parallel execution plans, each representing a unique way to parallelize the model by combining various parallelisms (i.e., DP, PP, TP). APEX then evaluates each plan by estimating the execution time of serving the input requests through simulation. APEX performs dynamism-aware simulation that is capable of modeling the characteristics of iteration-level batching. After simulation, APEX provides a comprehensive evaluation for each execution plan with key LLM serving metrics, such as time per output token (TPOT), time to first token (TTFT), and energy consumption, among others (Miao et al., 2023). This detailed evaluation enables service providers to select execution plans optimized for various objectives (e.g., minimizing latency or energy consumption) while meeting service-level objectives (SLOs).
Despite performing complex simulations, APEX remains highly scalable by leveraging the repetitive structure of transformer models, significantly reducing design space and enabling simulations at trillion-scale models across multi-node clusters. To model state-of-the-art LLM serving systems, APEX supports a broad range of LLMs, parallelisms, quantization formats, and device clusters. Recognizing the rapid evolution of LLM serving, APEX is designed to be modular and extensible, allowing easy integration of new models and hardware with minimal effort. Additionally, APEX leverages operation-level profilings to estimate the system execution time, which allows APEX to adapt to different clusters by collecting performance data from the underlying platform. Note that APEX not only supports GPU clusters but also ASIC clusters such as TPU (Jouppi et al., 2023) and Intel Gaudi (Intel, 2024). We evaluate APEX using a myriad of LLMs, device clusters, and distinct workloads. APEX consistently identifies optimal execution plans that outperform heuristic plans while maintaining low simulation overhead. APEX is open-sourced: https://github.com/microsoft/apex_plus. Our key contributions are:
@@
  • Automated parallel execution: We propose APEX, a simulator that automatically finds an optimal parallel execution plan for a given LLM, a device cluster, and input requests.
  • Dynamism-aware simulation: APEX models iteration-level batching, capturing real-world LLM serving behavior.
  • Accurate & effective: APEX delivers accurate speedup predictions, with an average relative error of just 10.7%. APEX finds optimal plans up to 3.37$\times$ faster than heuristic plans, and energy-optimal plans that reduce energy consumption by up to 45% compared to latency-optimal plans.
  • Broad compatibility and extensibility: APEX supports various LLMs, parallelisms, quantizations, and device clusters, while allowing easy extension to new configurations.
  • Cost-effective & scalable: APEX identifies an optimal plan within 15 minutes on a CPU, making it 71$\times$ faster and 1234$\times$ more cost-effective than cloud-based GPU deployment. It also maintains similar simulation overhead when scaling from billion-scale to trillion-scale models.
@@
  • Comprehensive evaluation: APEX reports key LLM-serving metrics, which can help service providers meet SLOs.
@@ 2.1. LLM Inference
Large Language Model (LLM) inference takes a user prompt as input, and generates a response token by token. The sequential token generation process is also known as auto-regressive generation (Gregor et al., 2014), which utilizes previously generated tokens as input to predict the next token. The inference process of LLMs consists of two stages:
1. Prefill Stage: In the prefill stage, LLM processes the input request (i.e., prompt) to set up intermediate states (keys and values) that are used to predict the first token. Unlike token generation, the computation of prefill does not rely on previously generated output tokens, allowing the tokens of the input request to be processed in parallel all at once. This high degree of parallelism makes the prefill stage compute-bound (Zhong et al., 2024).
2. Decode Stage: During the decode stage, LLM generates output tokens autoregressively until a stopping criterion is met. The generation of each token depends on all previous tokens’ output states (keys and values). The generation phase is often memory-bound (Lin et al., 2024), as its latency is primarily determined by the data transfer time of reading the output states from memory.
2.2. Device Clusters and Interconnections
To achieve high serving throughput and accommodate the large model size, LLMs are often deployed on a cluster of devices. The devices are connected in a tree-based network topology, which is one of the most popular topologies for node deployment. The memory bandwidth and latency is uniform within each level. Figure 1 illustrates an example of a commonly used two-level device cluster. Devices within the same node are typically connected by high-bandwidth interconnects like NVLink (Li et al., 2020a) at level 1, while devices across different nodes are connected via inter-node networks such as InfiniBand (Liu and Guitart, 2022) at level 2. In addition to GPUs, ASIC clusters are also emerging, such as TPU and Gaudi clusters (Jouppi et al., 2023; Intel, 2024). These clusters also utilize tree-based network topologies for interconnection and can be abstracted similarly as GPU clusters (Choquette, 2023).
@@ 2.3. LLM Serving and Iteration-Level Batching
LLMs are often hosted by service providers in the cloud (Li et al., 2024). Users send requests to these providers through APIs or chatbots and then receive responses generated by the LLMs. For LLM serving, it is critical to achieve both low latency and high throughput: low latency is necessary to meet service-level objectives, while high throughput enables service providers to serve a large number of users simultaneously. Due to the auto-regressive nature of LLMs, the lengths of the generated responses can vary significantly. Consequently, the commonly used static batching leads to suboptimal performance in LLM serving, as all the batched requests need to wait for the longest response to be generated to proceed to batch new requests. To overcome this inefficiency, iteration-level batching (Yu et al., 2022) is proposed. Iteration-level batching continuously schedules newly arrived request(s) into the existing batch whenever GPU memory becomes available, rather than waiting for all the requests to be completed. Iteration-level batching significantly improves the serving throughput and is widely adopted in state-of-the-art LLM serving systems (Kwon et al., 2023; Agrawal et al., 2024b; Zhong et al., 2024; NVIDIA, 2024).
Figure 1. An example two-level device cluster. Memory bandwidth and latency are uniform within the same level.
@@ 2.4. LLM Parallelisms
Data Parallelism (DP): In DP (Li et al., 2020b), the model is replicated across multiple devices, and input requests are divided into mini-batches, each processed independently by a replica. This approach avoids communication overhead during execution, but comes at the cost of increased memory usage due to model replication, thereby reducing available memory for batching requests.
@@
Expert Parallelism (EP): EP (Rajbhandari et al., 2022) is a special type of parallelism for Mixture of Experts (MoE) models (Cai et al., 2024). EP deploys different experts across various devices. For MoE models, only a subset of experts are activated for each input token. Thus, EP can result in workload imbalance and resource under-utilization if none of the experts on the device are activated; while this can be avoided by adopting TP, where each device is assigned with a partition of all the experts, EP incurs a lower communication overhead than TP.
@@ 3.1. System Overview
We depict the workflow of APEX in Figure 2. User-provided input data is shown in yellow-colored boxes, and built-in system data is represented by red-colored cylinders. Given an LLM and a device cluster, APEX generates multiple parallel execution plans using built-in Parallel Templates. Each generated executon plan represents a distinct strategy for mapping the LLM across the cluster. The Batching Module takes these generated plans and a trace of requests as inputs, dynamically generating the processing batch based on request arrival time and resource constraints. Specifically, the Batching Module determines, at each iteration, whether a request should be added to or removed from the processing batch and reports the active requests to the LLM Serving Simulator; the Simulator then uses profiling results to estimate execution time and energy consumption for processing the active requests, while also tracking system metrics such as time to first token (TTFT), time per output token (TPOT), P95 latency, among others. Once all requests are processed, the simulator generates a simulation report for each parallel execution plan. APEX can optimize towards different objectives by selecting the most suitable execution plan based on a parametrizable target metric: It can optimize for energy-efficiency by minimizing energy per inference, or can prioritize throughput by reducing end-to-end latency, enabling flexible adaptation to diverse system goals. Suppose that we aim to find the latency-optimal plan for serving Llama-3.1 70B on a cluster with 8 Nvidia H100 GPUs. APEX first generates multiple execution plans, such as 8-way tensor parallelism, or a 4-way tensor parallelism combined with 2-way pipeline parallelism. Next, it utilizes a trace of requests as input for simulation, such as the Azure LLM Traces (Patel et al., 2024). APEX simulates iteration-level batching to organize arriving requests into the processing batch. Finally, the LLM Serving Simulator estimates the execution time of each plan and select the one with the lowest end-to-end latency for serving Llama-3.1 70B on the cluster.
Figure 3. The Parallel Execution Plan Generator produces plans that map a given LLM onto the target cluster using distinct parallelization strategies
3.2. Parallel Execution Plan Generator
In this section, we describe the details of the Parallel Execution Plan Generator, which constructs various parallel execution plans to map the input LLM onto the target device cluster. The workflow of the Plan Generator is illustrated in Figure 3. The process begins by converting the input LLM into Transformer IR, an intermediate representation that provides a canonical abstraction for transformer-based models. Using this representation, the Plan Generator produces multiple parallel schemes based on the Parallel Templates. A parallel scheme maps the model onto a logical device cluster (Zhao et al., 2020), a virtual overlay that abstracts hardware details without network topology information. Next, the Device Mapper assigns logical devices to physical devices, generating executable parallel execution plans, which serve as inputs for the Batching Module. This two-stage mapping approach (first to the logical and then to physical device cluster) reduces the design complexity. We discuss each module in detail below.
3.2.1. IR Converter and Transformer IR
The Plan Generator converts an LLM into Transformer IR, a canonical representation of transformer-based models, as illustrated in Figure 4. We develop an IR converter that parses an LLM’s configuration file, extracting parameters such as the number of attention heads and layers, and translates the model into Transformer IR. In Transformer IR, an LLM is represented as a chain of cells, where each cell corresponds to key transformer operations, such as multi-head attention or the feedforward network. For example, GPT-3 (Brown et al., 2020) can be represented as a sequence of multi-head attention (MHA) cells and multi-layer perceptron (MLP) cells, while Llama-3.1 (Dubey et al., 2024) is modeled as a chain of group query attention (GQA) cells (Ainslie et al., 2023) and SwiGLU cells (Shazeer, 2020). Each cell consists of multiple tasks, such as individual attention heads in an MHA cell or experts in a mixture-of-experts (MoE) cell. Given that LLMs typically consist of repeated structures, we define the smallest set of nonrepetitive adjacent cells as a block. Using Transformer IR, an LLM is uniformly represented as multiple identical block, where each block consists of cells, and each cell contains multiple tasks. Transformer IR abstracts an LLM by focusing solely on key operations like MHA and MLP, while omitting operations like tokenization and position embeddings, which are less relevant for model parallelization. By focusing on key operations and leveraging the repetitive nature of LLMs through blocks, Transformer IR reduces both the search space and the simulation runtime for APEX. Furthermore, Transformer IR serves as a foundation for developing our Parallel Templates, which we describe in the following subsection.